OpenStudy (carlyleukhardt):

How many cups are in 10 quarts? A.25 B.30 C.36 D.40

OpenStudy (aaronandyson):

1 cup = 0.25 quartz x cups = 10 quartz Cross multiple 10*1 = 0.25*x Isolating x 10/0.25 = x Removing the decimal 10/25/100 = x 10*100/25 = 25 times 4 = 100 10*4 = x 40 = x Therefore D.